# OUJ-FTC-24

### v4

@@ 1 Introduction
In this work we explore whether mixing between an additional elementary field $\Psi$, coupled to the top quark via a new Yukawa interaction, and the toponium bound state can change the intrinsic width of toponium and account for the observed excess. Namely, the width $\Gamma_{\eta_{t}}\!\sim\!7~\mathrm{GeV}$ fiitted by the CMS[9] is larger than the nominal expectation for toponium, $\Gamma_{\eta_{t}}\!\sim\!3~\mathrm{GeV}$[13]. We interpret this discrepancy as arising from mixing with an additional elementary state. Moreover, even if future fittings prefer a value different from $7~\mathrm{GeV}$, such mixing can still account for it—for example, it can reduce the intrinsic width below $3~\mathrm{GeV}$, thereby lengthening the lifetime of the mixed mass eigenstate. As another motivation, [14] suggests that if only the bound state $\eta_{t}$ exists within SM, the stability of the Higgs vacuum may become unstable. Therefore, we also intend to introduce an additional elementary field to resolve this issue. When the bound state $\eta_{t}$ mixes with $\Psi$, the mass eigenstates $\Psi^{\prime}$ and $\eta_{t}^{\prime}$ are formed. We propose and analyze such mixing scenarios and delineate the phenomenologically viable parameter space, with particular emphasis on the magnitude of the mixing angle.
Renormalization-group (RG) control is essential even in the presence of the composite state $\eta_{t}$. To handle this, we adopt the Bardeen–Hill–Lindner (BHL) framework[14, 15, 16], in which the composite bound state is treated as an effective elementary field while imposing the compositeness condition $Z_{2}(\mu=\Lambda)=0$ at a cutoff scale $\Lambda$.
@@ 3 Minimal scenario to explain the excess in LHC
To include the non-perturbative formation of toponium, we follow the Bardeen-Hill-Lindner (BHL) prescription [15, 16, 14]. At a cutoff scale $\Lambda$, an effective four-fermion interaction for top quarks induces a wavefunction renormalization for $\eta_{t}$, and we impose the boundary condition $Z_{2}(\Lambda)=0$. This allows us to treat $\eta_{t}$ as an effective elementary field for $\mu\leq\Lambda$ and provides a consistent low-energy Effective Field Theory (EFT). In this framework we set the vacuum expectation values (VEVs) of $\Psi$ and $\eta_{t}$ to zero.
The resulting Yukawa sector captures effects beyond the perturbative regime and governs the top-quark dynamics. In addition to $y_{\eta_{t}}$ (between $\eta_{t}$ and the top), we also include interactions involving the new field $\Psi$. The additional Yukawa terms and the tree-level scalar potential are given by:
@@
Moreover, $\eta_{t}$ and $\Psi$ are constructed as doublets that, in analogy to the Higgs field, do not acquire VEV, thereby preserving gauge invariance. Their neutral components are CP-even; however, since the off-diagonal entries of the mass-mixing matrices with each of the Higgs fields vanish, they do not mix. Consequently, the constraints from scalar mixing angles at the LHC can be safely neglected.
@@
We search for the parameter $y_{\Psi}(\mu)$ that satisfies MPP conditions. For this reason, we examine the one-loop RGEs [Appendix A]. Also, we simply put $h=\mu$ since the effect on the effective potential is negligibly small and $h(\mu)$ is the renormalized running field. [^2] We do not distinguish between the bare field and the renormalized running field as its wave function renormalization is so small.. We use [28, 29]. Using the top quark mass $M_{t}=172.69\pm 0.30\ \mathrm{GeV}$ and the strong gauge coupling constant $\alpha_{s}(M_{Z})=0.1179\pm 0.0010$[29], we numerically calculate the parameter that approximately satisfies the MPP conditions. In this case, $y_{\Psi}(\mu)$ indirectly affects the effective potential through the RGEs. The MPP scale is found to be $\mu_{c}\lesssim 10^{12.3}\ \mathrm{GeV}$. To suppress potential corrections from additional new particles, we set the upper limit to $\mu_{c}=10^{12.3}\,\mathrm{GeV}$. In short, the MPP shift is minimized and unknown heavy states effects are diluted under standard EFT matching. We also take $\Lambda=370$$400\,\mathrm{GeV}$, since the effect persists up to several tens of GeV above the toponium mass[30]. Moreover, since the lower bound lies around $345~\mathrm{GeV}$[30], we assume the four-fermion interaction emerges near this scale. Accordingly, the terms involving $\eta_{t}$ are run between $345~\mathrm{GeV}$ and the cutoff $\Lambda$. For $\Lambda>\mu$, such a four-fermion interaction is absent, since this coupling is intended only to describe the near-threshold region. Consequently, it does not conflict with the relevant experimental constraints. The results are as follows:
@@
And the mixing angle $\theta$ between $\eta_{t}$ and $\Psi$ is likewise essentially determined. Note that we are not dealing with stable particles here, so the mixing angle can be written as a complex number, $\theta=\theta_{R}+i\theta_{C}$. As will be discussed later, since we assume the decay widths are sufficiently small, we may approximate $|\theta|\sim|\theta_{R}|$; hence the allowed range is $|\theta|\leq 45^{\circ}$, and it is determined as follows:
@@ 4 In case of the 2HDM to explain the excess in LHC
It should be noted that the so-called oblique parameters ($S$, $T$, and $U$) must be taken into account between the 2HDM and $\eta_{t}$[26]. After the basis transformation in the 2HDM, mixing arises between each mass component (the new scalar, pseudo-scalar, and charged scalar) and the corresponding components of $\eta_{t}$. However, if we assume that the mass of toponium remains approximately $345$ GeV even after the mixing, the constraints from the oblique parameters can be neglected, since in such a case the masses of each component on the 2HDM side also change only negligibly. Therefore, it is sufficient to consider only the deviations of the oblique parameters originating from the 2HDM.
@@ 5 Summary and Discussion
Motivated by the excess of $t\bar{t}$ events near threshold reported by the LHC experiments, we studied a framework in which a bound state of top quarks (toponium, $\eta_{t}$) mixes with an additional elementary field $\Psi$. Our analysis was carried out in an effective description that incorporates nonperturbative formation effects via the Bardeen-Hill-Lindner (BHL) prescription, supplemented with the Multicritical Point Principle (MPP) as a high-scale constraint on couplings.

### v5

@@ 1 Introduction
In this work, we consider mixing between an additional elementary field $\Psi$, which couples to the top quark through a new Yukawa interaction, and the toponium bound state. As a motivation, [13] suggests that if only the bound state $\eta_{t}$ exists within the SM, the stability of the Higgs vacuum may be driven in a more unstable direction, which can in some cases further destabilize the vacuum. Therefore, we also intend to introduce an additional elementary field to address this issue. When the bound state $\eta_{t}$ mixes with $\Psi$, the mass eigenstates $\Psi^{\prime}$ and $\eta_{t}^{\prime}$ are formed. We propose and analyze such mixing scenarios and delineate the phenomenologically viable parameter space, with particular emphasis on the magnitude of the mixing angle.
Renormalization Group control is essential even in the presence of the composite state $\eta_{t}$. To handle this, we adopt the Bardeen-Hill-Lindner (BHL) framework[13, 14, 15], in which the composite bound state is treated as an effective elementary field while imposing the compositeness condition $Z_{2}(\mu=\Lambda)=0$ at a cutoff scale $\Lambda$.
@@ 3 Minimal scenario to explain the excess in LHC
To include the short-distance effects near the threshold around toponium, we follow the Bardeen-Hill-Lindner (BHL) prescription[13, 14, 15]. At a cutoff scale $\Lambda$, an effective four-fermion interaction for top quarks induces a wavefunction renormalization for $\eta_{t}$, and we impose the boundary condition $Z_{2}(\Lambda)=0$. This allows us to treat $\eta_{t}$ as an effective elementary field for $\mu\leq\Lambda$ and provides a consistent low-energy Effective Field Theory (EFT). In this framework we set the vacuum expectation values (VEVs) of $\Psi$ and $\eta_{t}$ to zero.
The resulting Yukawa sector captures effects beyond the perturbative regime and governs the top-quark dynamics. In addition to $y_{\eta_{t}}$ (between $\eta_{t}$ and the top), we also include interactions involving the new field $\Psi$. The additional Yukawa terms and the additional tree-level scalar potential are given by:
@@
Moreover, $\eta_{t}$ and $\Psi$ are constructed as doublets that, in analogy to the Higgs field, do not acquire VEV, thereby preserving gauge invariance. Also, for convenience in writing a gauge-invariant EFT, we extend $\eta_{t}$ by introducing a neutral scalar component and a charged component. In this work, these additional components contribute effectively only at the loop-level, and we therefore consider their effects only to that extent in what follows. Accordingly, we do not address whether the scalar or charged $\eta_{t}$ components could be experimentally observable.
@@
We search for the parameter $y_{\Psi}(\mu)$ that satisfies MPP conditions. For this reason, we examine the one-loop RGEs [Appendix A]. Also, we simply put $h=\mu$ since the effect on the effective potential is negligibly small and $h(\mu)$ is the renormalized running field. [^2] We do not distinguish between the bare field and the renormalized running field as its wave function renormalization is so small.. We use [27, 28]. Using the top quark mass $M_{t}=172.69\pm 0.30\ \mathrm{GeV}$ and the strong gauge coupling constant $\alpha_{s}(M_{Z})=0.1179\pm 0.0010$[28], we numerically calculate the parameter that approximately satisfies the MPP conditions. In this case, $y_{\Psi}(\mu)$ indirectly affects the effective potential through the RGEs. The MPP scale is found to be $\mu_{c}\lesssim 10^{12.3}\ \mathrm{GeV}$. To suppress potential corrections from additional new particles, we set the upper limit to $\mu_{c}=10^{12.3}\,\mathrm{GeV}$. In short, the MPP shift is minimized and unknown heavy states effects are diluted under standard EFT matching. Also, since the threshold effects persist up to several tens of GeV above the toponium mass[29], we conservatively assume that the short-distance effects near the threshold extend over a similar energy range, and take their upper limit to be $\Lambda=370$$400,\mathrm{GeV}$. Moreover, since the lower bound lies around $345~\mathrm{GeV}$[29], we assume the four-fermion interaction emerges near this scale. Accordingly, the terms involving $\eta_{t}$ are run between $345~\mathrm{GeV}$ and the cutoff $\Lambda$. For $\Lambda>\mu$, such a four-fermion interaction is absent, since this coupling is intended only to describe the near-threshold region. Consequently, it does not conflict with the relevant experimental constraints. The results are as follows:
@@
And the mixing angle $\theta$ between $\eta_{t}$ and $\Psi$ (both being pseudo-scalar components) is likewise essentially determined. Note that we are not dealing with stable particles here, so the mixing angle can be written as a complex number, $\theta=\theta_{R}+i\theta_{C}$. As will be discussed later, since we assume the decay widths are sufficiently small, we may approximate $|\theta|\sim|\theta_{R}|$; hence the allowed range is $|\theta|\leq 45^{\circ}$, and it is determined as follows:
@@ 4 In case of the 2HDM to explain the excess in LHC
It should be noted that the so-called oblique parameters ($S$, $T$, and $U$) must be taken into account between the 2HDM and $\eta_{t}$[25]. After the basis transformation in the 2HDM, mixing arises between mass component (the new pseudo-scalar) and the corresponding component (pseudo-scalar) of $\eta_{t}$. However, if we assume that the mass of toponium remains approximately $345$ GeV even after the mixing, the constraints from the oblique parameters can be neglected, since in such a case the mass of component (the new pseudo-scalar) on the 2HDM side also change only negligibly. Therefore, it is sufficient to consider only the deviations of the oblique parameters originating from the 2HDM.
@@ 5 Summary and Discussion
Motivated by the excess of $t\bar{t}$ events near threshold reported by the LHC experiments, we studied a framework in which a bound state of top quarks (toponium, $\eta_{t}$) mixes with an additional elementary field $\Psi$. Our analysis was carried out in an effective description that incorporates short-distance effects near the threshold via the Bardeen-Hill-Lindner (BHL) prescription, supplemented with the Multicritical Point Principle (MPP) as a high-scale constraint on couplings.